# Multi-Order Hyperbolic Graph Convolution and Aggregated Attention for Social Event Detection

### v1

@@ 4 Methodology
This section describes our methodology, beginning with an overview of multi-order graph convolution and aggregated in hyperbolic space (Section 4.1). We then apply this method to unsupervised learning event detection (Section 4.2) and validate its effectiveness in supervised learning settings (Section 4.3).
@@ 4.1.1 Overview
Figure 5: The overall framework of this study: First, the data representation $\mathrm{X}_{i}$ and its corresponding adjacency matrix are obtained. Next, $\mathrm{X}_{i}$ is mapped into hyperbolic space via the $\exp_{o}(\cdot)$ function to produce its hyperbolic representation. Subsequently, by applying the $\log_{o}(\cdot)$ function, $\mathrm{X}_{i}$ is mapped onto its tangent space, yielding the representation $x_{i}$. The multi-order graph convolution network is then employed to derive the multi-order representation $h_{i}^{k}$. Building upon this, an attention-based network is used to generate the hyperbolic multi-order graph representation $h_{i}$. Finally, the convolution attention representation $h_{i}$ is mapped into a new hyperbolic space by means of the $\log_{o}(\cdot)$ function, resulting in the hyperbolic representation $\mathcal{H}_{i}$.
To perform multi-order graph convolution aggregation in hyperbolic space, our method follows four key steps: First, we initialize the node features from Euclidean space into hyperbolic space and map these features to the tangent space at the origin. Second, we perform convolution operations of different orders within the tangent space to obtain high-order information about these features. Third, we aggregate the convolution representations from various orders using an attention mechanism. This ensures that the most relevant features are emphasized during the aggregation process. Finally, the aggregated representation in the tangent space is mapped back to a new hyperbolic space using the exponential map. This step obtains in the final output a node feature representation of a high-order convolution aggregation of hyperbolic space. This structured approach combines the advantages of hyperbolic geometry and graph convolution, enabling efficient and meaningful feature learning.
@@ 4.1.2 Initial Hyperbolic Layer and Tangent Space Transformation
To initialize node features in hyperbolic space, we project Euclidean node features using the exponential map (Equation 1 and Equation 2 in Section 3.3). These features are then mapped to the tangent space at $o$ via the logarithmic map, enabling compatibility with Euclidean operations for graph neural network (GNN) processing.
4.1.3 Multi-Order Graph Attention Network
We capture multi-order relationships through adjacency matrices for each node $A$, consisting of three submatrices of dimension $n\times n$: $A_{\mathrm{along}}$, $A_{\mathrm{rev}}$, and $A_{\mathrm{loop}}$.
@@ 4.1.3 Multi-Order Graph Attention Network
The aggregated $K$-order adjacency matrix in hyperbolic space is defined as:
$A_{\mathrm{multi}}^{k}=\bigoplus_{\mathrm{subg}\in\{\mathrm{along,rev,loop}\}}A_{\mathrm{subg}}^{k}$ (5)
where $A_{\mathrm{subg}}^{k}$ represents $k$-order relationships from each subgraph $\mathrm{subg}$.
@@
where $f(\cdot)$ is the graph convolution function in hyperbolic space, and $\oplus$ denotes element-wise addition. The convolution function $f(p_{i},a^{k})$ is defined as:
@@ 4.1.4 Attention-Based Aggregation
$v_{i}^{k}=\mathrm{softmax}(s_{i}^{k})=\frac{\exp(V\tanh(Wx_{i}))}{\sum_{n=1}^{N}\exp(V\tanh(Wx_{j}))}$ (10)
Here, $W$ and $V$ are linear transformations that project node features to produce scalar attention scores, emphasizing important orders for each node’s representation in hyperbolic space.
@@ 4.1.5 Mapping from Tangent Space to Hyperbolic Space
4.2 An Unsupervised Multi-Order Hyperbolic Graph Convolution and Aggregated Attention for Social Event Detection
@@ 4.2.1 Overall Framework
Figure 6: The framework of the ’UMOHGCAASED’ model. $\mathbf{X}$ represents the obtained node feature, $\mathbf{X^{\prime}}$ signifies the augmented node feature of $\mathbf{X}$, $\mathcal{H}$ and $\mathcal{H^{\prime}}$ indicate the hyperbolic feature following hyperbolic multi-order aggregation, $\mathcal{E}$ and $\mathcal{E^{\prime}}$ denote the Euclidean spatial feature, and $\mathcal{Y}$ and $\mathcal{Y^{\prime}}$ are the ultimate feature representations.
@@ 4.2.2 Graph Data Augmentation
Graph data augmentation is essential for contrastive learning, especially in domains like computer vision (CV) and natural language processing (NLP). In the preliminary sections, we introduced the construction of graphs with nodes and relationships. Here, we use graph augmentation techniques to assess how our method affects node representation in hyperbolic space. Specifically, we select feature corruption as the primary augmentation technique for node features. Given a graph $\mathcal{G}=(\mathcal{N,A})$ and its augmented version $\mathcal{G^{\prime}}=(\mathcal{N^{\prime},A^{\prime}})$, the node set changes ($\mathcal{N}\neq\mathcal{N^{\prime}}$) while the adjacency matrix remains the same ($\mathcal{A}=\mathcal{A^{\prime}}$).
@@ 4.2.3 Multi-order Hyperbolic Space Node Feature Graph Convolution and Aggregation Encoding
$h_{i}^{l,\mathcal{H}}=\left(W^{l}\otimes c^{e-1}x_{i}^{l-1,\mathcal{H}}\right)\oplus c^{e-1}b^{l},$ (11)
where $W^{l}\otimes c^{e-1}x_{i}^{l-1,\mathcal{H}}$ represents the hyperbolic tangent space operation described in Equation 1:
$W\otimes x^{B}:=\exp_{o}^{c}(W\log_{o}^{c}(x^{B})),$ (12)
@@
$\mathbf{x}^{\mathcal{H}}\oplus_{c}^{\mathcal{H}}\mathbf{b}^{\mathcal{H}}=\exp_{\mathbf{x}^{\mathcal{H}}}^{c}\left(P_{o\to\mathbf{x}^{\mathcal{H}}}^{c}\left(\log_{o}^{c}(\mathbf{b}^{\mathcal{H}})\right)\right).$ (14)
By using the method in Section 4.1.3, we calculate the attention-weighted representation:
@@
and apple the method in Section 4.1.4:
@@ 4.2.4 Contrastive Loss
$\mathcal{E}=\log^{c}_{o}\left(\mathcal{H}\right),$ (19)
@@
$\mathcal{E}^{\prime}=\log^{c}_{o}\left(\mathcal{H}^{\prime}\right).$ (20)
@@
$\mathcal{L}_{MOUHGA}=\frac{1}{N+M}\left(\sum_{i=1}^{N}\mathcal{E}\left[\log\mathcal{D}\left(e_{i},\mathcal{Y}\right)\right]+\sum_{j=1}^{M}\mathcal{E}^{\prime}\left[\log\left(1-\mathcal{D}\left(e^{\prime}_{j},\mathcal{Y}\right)\right)\right]\right),$ (21)
where $\mathcal{Y}=\mathcal{R}(\mathcal{E})$ is derived using a readout function $\mathcal{R}(\cdot)$ defined as:
$\mathcal{R}(\mathcal{E})=\sigma\left(\frac{1}{N}\sum_{i=1}^{N}e_{i}\right),$ (22)
where $e_{i}\in\mathcal{E}$. Finally, a discriminator $\mathcal{D}(\cdot)$ Oord et al. (2018) maximizes the mutual information, formulated as:
$\mathcal{D}(e_{i},\mathcal{Y})=\rho\left(e_{i}\mathcal{W}\mathcal{Y}\right),$ (23)
where $\mathcal{W}$ is a learnable scoring matrix, and $\rho$ denotes the nonlinear logistic sigmoid function.
4.3 A Supervised Multi-Order Hyperbolic Graph Convolution and Aggregated Model for Social Event Detection
@@ 4.3.1 Overall Framework
Figure 7: The framework of the ’MOHGCAASED’ model. $\mathbf{X}$ represents the node characteristics, $\mathcal{H}$ signifies hyperbolic features subsequent to hyperbolic multi aggregation, $\mathcal{E}$ indicates Euclidean space features following the logarithmic transformation, and $\mathcal{Y}$ constitutes the ultimate feature representation.
@@ 4.3.2 Multi-Order Graph Convolution and Aggregation in Hyperbolic Space and Decoder Design
$\mathcal{E}=\log^{c}_{o}\left(\mathcal{H}\right).$ (25)
@@
$\mathcal{Y}=\mathbf{W}\cdot\mathcal{E}+\mathbf{b}.$ (26)
@@ 5.1 Datasets
The datasets used in this experiment include: Our model primarily targets social media data, so we selected the real-world Twitter McMinn et al. (2013) dataset as our primary dataset. However, since the Twitter dataset is a large-scale dataset, contrastive learning with it may lead to out-of-memory errors. To address this, we used a balanced, smaller “mini-Twitter” dataset derived from the original dataset. Additionally, we used the Cora Sen et al. (2008) dataset, and the Citeseer Sen et al. (2008) dataset to evaluate the effectiveness of modeling in hyperbolic space. Table 2 provides an overview of these datasets.
@@
Dataset Num. of Classes Num. of Nodes Num. of Features
@@ 5.2 MOHGCAA Experiments in Unsupervised Settings
In this subsection, we evaluate the MOHGCAA model under unsupervised settings. Section 5.2.1 introduces the baseline experiments, and Section 5.2.2 provides a detailed description of the experimental environment and parameter settings. The remainder of this section addresses questions Q1 and Q2 with respect to the MOHGCA model.
@@ 5.2.1 Baseline Models
  • DGI Veličković et al. : A single-branch graph contrastive learning model. It generates negative samples by either augmenting data or corrupting the original graph structure, excelling in unsupervised learning of node representations through mutual information maximization.
  • GraphCL You et al. : A dual-branch graph contrastive learning framework. It performs augmentations on the input graph to generate two distinct views, capturing structural and semantic similarities to enhance representation learning.
  • GCN Kipf and Welling (b): A foundational graph neural network model leveraging spectral graph theory to generalize convolutional operations to graph structures. By utilizing graph Laplacians, GCN performs layer-wise aggregation of features from local neighborhoods, enabling efficient semi-supervised learning.
  • HGCN Chami et al. (b): An extension of GCN to hyperbolic space, designed for hierarchical and complex graph data. HGCN capitalizes on the unique properties of hyperbolic geometry to represent hierarchical structures with exponential capacity, achieving superior results in tasks like node classification and link prediction.
  • HNN Ganea et al. : A hyperbolic neural network that operates directly in hyperbolic space to model hierarchical and structured datasets. By redefining core neural operations in hyperbolic geometry, HNN provides enhanced performance in classification and representation learning of complex data.
  • HyboNet Chen et al. : A sophisticated hyperbolic neural framework fully utilizing the Lorentz model of hyperbolic geometry. HyboNet avoids inefficient mappings between Euclidean and hyperbolic spaces, incorporating hyperbolic rotations and attention mechanisms to achieve compact, powerful embeddings for relational and hierarchical data.
  • UHSED Qiu et al. (a): An unsupervised hyperbolic graph-based model for social event detection, specifically tailored for heterogeneous and unlabelled social media datasets. By constructing a unified social message graph and employing contrastive learning, UHSED captures both semantic and structural information effectively.
@@ 5.2.2 Parameter Settings
All the experiments with the MOHGCA model were conducted on an NVIDIA V100 GPU with 12 24-core Intel Xeon processors.
@@ 5.2.3 Evaluation Metrics
Since our proposed method is an unsupervised learning model, we adopted Micro-F1 and Macro-F1 scores to evaluate the accuracy of the detection results, following the methodology of prior studies Veličković et al. ; You et al. . These metrics were chosen for their robustness in assessing performance across imbalanced datasets, with Micro-F1 capturing overall classification performance and Macro-F1 emphasizing class-specific performance.
@@ 5.3.1 Baseline Models
  • Word2vec Mikolov et al. :An approach employing universal embedding representation for the social event identification.
  • LDA Blei et al. : A traditional statistical approach is employed for the social event detection.
  • WMD Kusner et al. : A similarity-based measurement method for detecting social events by computing the similarity between messages.
  • BERT Devlin : A robust language representation model that plays a crucial role in numerous state-of-the-art social event detection models.
  • KPGNN Cao et al. (b): A social event detection model based on heterogeneous information networks (HINs), primarily designed for incremental social event detection. Its offline performance slightly outperforms that of PP-GCN.
  • FinEvent Peng et al. (b): A social event detection model that leverages incremental and cross-lingual social messages, employing reinforcement learning to enhance detection performance.
  • HSED Qiu et al. (a): A hyperbolic space-based social event detection model.
  • HNN Ganea et al. : Another hyperbolic neural network model applied for social event detection.

### v2

@@ 4 Methodology
This section describes our methodology, beginning with an overview of multi-order graph convolution and aggregation in hyperbolic space (Section 4.1). We then apply this method to unsupervised learning event detection (Section 4.2) and validate its effectiveness in supervised learning settings (Section 4.3).
@@ 4.1.1 Overview
Figure 5: The overall framework of this study: First, the data representation $\mathrm{X}_{i}$ and its corresponding adjacency matrix are obtained. Next, $\mathrm{X}_{i}$ is mapped into hyperbolic space via the $\exp_{o}(\cdot)$ function, producing its hyperbolic representation. Then, the $\log_{o}(\cdot)$ function is applied to project $\mathrm{X}_{i}$ onto its tangent space, yielding the representation $x_{i}$.On this basis, a multi-order graph convolution network is employed to derive the multi-order representation $h_{i}^{k}$. Subsequently, an attention-based network is used to generate the hyperbolic multi-order graph representation $h_{i}$. Finally, the convolution attention representation $h_{i}$ is mapped into a new hyperbolic space through the $\log_{o}(\cdot)$ function, resulting in the hyperbolic representation $\mathcal{H}_{i}$.
To perform multi-order graph convolution aggregation in hyperbolic space, our method follows four key steps: First, we initialize the node features from Euclidean space into hyperbolic space and map these features to the tangent space at the origin. Second, we perform convolution operations of different orders within the tangent space to obtain high-order information about these features. Third, we aggregate the convolution representations from various orders using an attention mechanism. This ensures that the most relevant features are emphasized during the aggregation process. Finally, the aggregated representation in the tangent space is mapped back to a new hyperbolic space using the exponential map. This step obtains in the final output a node feature representation of a high-order convolution aggregation of hyperbolic space. This structured approach combines the advantages of hyperbolic geometry and graph convolution, enabling efficient and meaningful feature learning. The overall framework can refer the Figure5.
@@ 4.1.2 Initial Hyperbolic Layer and Tangent Space Transformation
To initialize node features in hyperbolic space, we project Euclidean node features using the exponential map (Equation 1 and Equation 2 in Section 3.3). These features are then mapped to the tangent space at $o$ via the logarithmic map (Equation 3 and Equation 4 in Section 3.3), enabling compatibility with Euclidean operations for graph neural network (GNN) processing.
4.1.3 Multi-Order Graph Convolution Network
We capture multi-order relationships through adjacency matrices $A$ for each node, consisting of three submatrices of dimension $n\times n$: $A_{\mathrm{along}}$, $A_{\mathrm{rev}}$, and $A_{\mathrm{loop}}$.
@@ 4.1.3 Multi-Order Graph Convolution Network
where $f(\cdot)$ is the graph convolution function in hyperbolic space, $k$ is the orders to aggregate, and $\oplus$ denotes element-wise addition. The convolution function $f(p_{i},a^{k})$ is defined as:
@@ 4.1.4 Attention-Based Aggregation
$v_{i}^{k}=\mathrm{softmax}(s_{i}^{k})=\frac{\exp(W_{V}\tanh(Wx_{i}))}{\sum_{n=1}^{N}\exp(W_{V}\tanh(Wx_{j}))}$ (9)
. Here, $W$ and $W_{V}$ serve as linear transformations that project node features into scalar attention scores, thereby highlighting the most relevant orders for each node’s representation in hyperbolic space.
@@ 4.1.5 Mapping from Tangent Space to Hyperbolic Space
4.2 Multi-Order Unsupervised Hyperbolic Graph Convolution and Aggregated Attention for Social Event Detection
@@ 4.2.1 Overall Framework
Figure 6: The framework of the Multi-Order Unsupervised Hyperbolic Graph Convolution and Aggregated Attention for Social Event Detection (MOUHGCAASED) model. $\mathbf{X}$ represents the obtained node feature, $\mathbf{X^{\prime}}$ signifies the augmented node feature of $\mathbf{X}$, $\mathcal{H}$ and $\mathcal{H^{\prime}}$ indicate the hyperbolic feature following hyperbolic multi-order aggregation, $\mathcal{X}$ and $\mathcal{X^{\prime}}$ denote the Euclidean spatial feature, and $\mathcal{Y}$ and $\mathcal{Y^{\prime}}$ are the ultimate feature representations.
@@ 4.2.2 Graph Data Augmentation
Graph data augmentation is essential for contrastive learning, especially in domains like computer vision (CV) and natural language processing (NLP). In the preliminary sections, we introduced the construction of graphs with nodes and relationships. Here, we use graph augmentation techniques to assess how our method affects node representation in hyperbolic space. Specifically, we select feature corruption as the primary augmentation technique for node features. Given a graph $\mathcal{G}=(V,\mathcal{A})$ and its augmented version $\mathcal{G^{\prime}}=(V^{\prime},\mathcal{A^{\prime}})$, the node set changes ($V\neq V^{\prime}$) while the adjacency matrix remains the same ($\mathcal{A}=\mathcal{A^{\prime}}$).
@@ 4.2.3 Multi-order Hyperbolic Space Node Feature Graph Convolution and Aggregation Encoding
$h_{i}^{l,\mathcal{H}}=\left(W^{l}\otimes_{c}^{\mathcal{H}}x_{i}^{l-1,\mathcal{H}}\right)\oplus_{c}^{\mathcal{H}}b^{l},$ (10)
where $W^{l}\otimes_{c}^{\mathcal{H}}x_{i}^{l-1,\mathcal{H}}$ represents the hyperbolic tangent space operation described in Equation 1:
$W\otimes x^{\mathcal{B}}:=\exp_{o}^{c}(W\log_{o}^{c}(x^{B})),$ (11)
@@
$\mathbf{x}^{\mathcal{H}}\oplus_{c}^{\mathcal{H}}\mathbf{b}^{\mathcal{H}}=\exp_{\mathbf{x}^{\mathcal{H}}}^{c}\left(P_{o\to\mathbf{x}^{\mathcal{H}}}^{c}\left(\log_{o}^{c}(\mathbf{b}^{\mathcal{H}})\right)\right)$ (13)
, where $\mathbf{b}^{\mathcal{H}}$ is the bias in hyperbolic space $\mathcal{H}_{c}^{n}$ with curvature $c$ and the parallel transport $P_{o\to\mathbf{x}^{\mathcal{H}}}^{c}$ is the function from the origin $o$ to the hyperbolic point $\mathbf{x}^{\mathcal{H}}$ under curvature $c$. By using the method in Section 4.1.3, we calculate the attention-weighted representation:
@@
and apply the method in Section 4.1.4:
@@ 4.2.4 Contrastive Loss
$\mathcal{X}=\log^{c}_{o}\left(\mathcal{H}\right),$ (18)
@@
$\mathcal{X}^{\prime}=\log^{c}_{o}\left(\mathcal{H}^{\prime}\right).$ (19)
@@
$\mathcal{L}_{MOUHGA}=\frac{1}{N+M}\Biggl{(}\sum_{i=1}^{N}\mathcal{X}\bigl{[}\log\mathcal{D}\bigl{(}e_{i},\mathcal{Y}\bigr{)}\bigr{]}\\ +\sum_{j=1}^{M}\mathcal{X}^{\prime}\bigl{[}\log\bigl{(}1-\mathcal{D}(e^{\prime}_{j},\mathcal{Y})\bigr{)}\bigr{]}\Biggr{)}$ (20)
, where $\mathcal{Y}=\mathcal{R}(\mathcal{X})$ is derived using a readout function $\mathcal{R}(\cdot)$ defined as:
$\mathcal{R}(\mathcal{X})=\sigma\left(\frac{1}{N}\sum_{i=1}^{N}e_{i}\right),$ (21)
where $e_{i}\in\mathcal{X}$. Finally, a discriminator $\mathcal{D}(\cdot)$ [40] maximizes the mutual information, formulated as:
$\mathcal{D}(e_{i},\mathcal{Y})=\rho\left(e_{i}\mathcal{W}\mathcal{Y}\right)$ (22)
, where there $\mathcal{W}$ is a learnable scoring matrix, and $\rho$ denotes the nonlinear logistic sigmoid function.
4.3 Multi-Order Supervised Hyperbolic Graph Convolution and Aggregated Model for Social Event Detection
@@ 4.3.1 Overall Framework
Figure 7: The framework of the Multi-Order Hyperbolic Graph Convolution and Aggregated Model for Social Event Detection (MOHGCAASED) model. $\mathbf{X}$ represents the node characteristics, $\mathcal{H}$ signifies hyperbolic features subsequent to hyperbolic multi aggregation, $\mathcal{X}$ indicates Euclidean space features following the logarithmic transformation, and $\mathcal{Y}$ constitutes the ultimate feature representation.
@@ 4.3.2 Multi-Order Graph Convolution and Aggregation in Hyperbolic Space and Decoder Design
$\mathcal{X}=\log^{c}_{o}\left(\mathcal{H}\right).$ (24)
@@
$\mathcal{Y}=\mathbf{W}\cdot\mathcal{X}+\mathbf{b}.$ (25)
@@ 5.1 Datasets
The datasets used in this experiment include: Our model primarily targets social media data, so we selected the real-world Twitter [41] dataset as our primary dataset. However, since the Twitter dataset is a large-scale dataset, contrastive learning with it may lead to out-of-memory errors. To address this, we used a balanced, smaller “mini-Twitter” dataset derived from the original dataset. Additionally, we used the Cora and the Citeseer [42] datasets to evaluate the effectiveness of modeling in hyperbolic space. Table 2 provides an overview of these datasets.
@@
Dataset ##\## of Classes ##\## of Nodes ##\## of Features
@@ 5.2 MOHGCAA Experiments in Unsupervised Settings
In this subsection, we evaluate the MOHGCAA model under unsupervised settings. Section 5.2.1 introduces the baseline experiments, and Section 5.2.2 provides a detailed description of the experimental environment and parameter settings. The remainder of this section addresses questions Q1 and Q2 with respect to the Multi-Order Supervised Hyperbolic Graph Convolution and Aggregated Model for Social Event Detection (MOUHGCAASED) model.
@@ 5.2.1 Baseline Models
  • DGI [43]: A single-branch graph contrastive learning model. It generates negative samples by either augmenting data or corrupting the original graph structure, excelling in unsupervised learning of node representations through mutual information maximization.
  • GraphCL [44]: A dual-branch graph contrastive learning framework. It performs augmentations on the input graph to generate two distinct views, capturing structural and semantic similarities to enhance representation learning.
  • GCN [25]: A foundational graph neural network model leveraging spectral graph theory to generalize convolution operations to graph structures. By utilizing graph Laplacians, GCN performs layer-wise aggregation of features from local neighborhoods, enabling efficient semi-supervised learning.
  • HGCN [45]: An extension of GCN to hyperbolic space, designed for hierarchical and complex graph data. HGCN capitalizes on the unique properties of hyperbolic geometry to represent hierarchical structures with exponential capacity, achieving superior results in tasks like node classification and link prediction.
  • HNN [46]: A hyperbolic neural network that operates directly in hyperbolic space to model hierarchical and structured datasets. By redefining core neural operations in hyperbolic geometry, HNN provides enhanced performance in classification and representation learning of complex data.
  • HyboNet [47]: A sophisticated hyperbolic neural framework fully utilizing the Lorentz model of hyperbolic geometry. HyboNet avoids inefficient mappings between Euclidean and hyperbolic spaces, incorporating hyperbolic rotations and attention mechanisms to achieve compact, powerful embeddings for relational and hierarchical data.
  • UHSED [8]: An unsupervised hyperbolic graph-based model for social event detection, specifically tailored for heterogeneous and unlabelled social media datasets. By constructing a unified social message graph and employing contrastive learning, UHSED captures both semantic and structural information effectively.
@@ 5.2.2 Parameter Settings
All the experiments with the model were conducted on an NVIDIA GeForce RTX 3090 GPU with AMD EPYC 7302 64-core CPU processors. For other experimental hyperparameters, please refer to Table 3
@@ 5.2.3 Evaluation Metrics
Since our proposed method is an unsupervised learning model, we adopted Micro-F1 and Macro-F1 scores to evaluate the accuracy of the detection results, following the methodology of prior studies [43, 44]. These metrics were chosen for their robustness in assessing performance across imbalanced datasets, with Micro-F1 capturing overall classification performance and Macro-F1 emphasizing class-specific performance.
@@ 5.3.1 Baseline Models
  • Word2vec [11]:An approach employing universal embedding representation for the social event identification.
  • LDA [32]: A traditional statistical approach is employed for the social event detection.
  • WMD [48]: A similarity-based measurement method for detecting social events by computing the similarity between messages.
  • BERT [49]: A robust language representation model that plays a crucial role in numerous state-of-the-art social event detection models.
  • KPGNN [4]: A social event detection model based on heterogeneous information networks (HINs), primarily designed for incremental social event detection. Its offline performance slightly outperforms that of PP-GCN.
  • FinEvent [50]: A social event detection model that leverages incremental and cross-lingual social messages, employing reinforcement learning to enhance detection performance.
  • HSED [8]: A hyperbolic space-based social event detection model.
  • HNN [46]: Another hyperbolic neural network model applied for social event detection.